OpenStudy (mathmath333):

The difference between the radius of the front wheel and the rear wheel of a tractor is \(2\) feet . The ratio of the number of the revolutions made by the front and rear wheels of the tractor in a journey of \(13200\) feet is \(5:3\). How many more revolutions are made by the front wheel than rear wheel in the whole journey.

OpenStudy (mathmath333):

@campbell_st

OpenStudy (mathmath333):

this might be useful \(\large \color{black}{\begin{align} f-r=2\\~\\ \dfrac{m(2\pi f )}{n(2\pi r)}=\dfrac{5}{3}\\~\\ \end{align}}\)

OpenStudy (campbell_st):

well both wheels travel 13200 ft... let r be the radius of the rear so the rear revolutions are is 13200/2pir front revolutions 13200/2pi(r - 2) so then given the ratio 5 : 3 13200 13200 -------- : ------ 2pi(r -2) 2pi r there you go... thats a start

OpenStudy (mathmath333):

\(\large \color{black}{\begin{align} \dfrac{\dfrac{13200}{2\pi (r+2)}}{\dfrac{13200}{2\pi (r)}}=\dfrac{5}{3}\hspace{.33em}\\~\\ \end{align}}\) is this correct?

OpenStudy (mathmath333):

@Michele_Laino

OpenStudy (michele_laino):

I think that the problem can be modeled by the subsequent equations: \[\left\{ \begin{gathered} D = d + 4 \hfill \\ \frac{N}{n} = \frac{5}{3} \hfill \\ ND = nd \hfill \\ \pi DN = 13200 \hfill \\ \end{gathered} \right.\]

OpenStudy (michele_laino):

where D, N are the diameter and the frequency of the front wheel, whereas d, and n are the correspondent quantities of the rear wheel

OpenStudy (mathmath333):

frequency ?

OpenStudy (michele_laino):

number of revolutions per second

OpenStudy (michele_laino):

oops...where D, N are the diameter and the frequency of the rear wheel, whereas d, and n are the correspondent quantities of the front wheel

OpenStudy (mathmath333):

d and n means which quantities ?

OpenStudy (michele_laino):

d and n are the diameter and the frequency of the front wheel